OpenStudy (anonymous):

Check my answer Please!! A large emerald with a mass of 378.24 grams was recently discovered in a mine. If the density of the emerald is 2.76grams over centimeters cubed, what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point. I got 137.04

OpenStudy (perl):

$$ \Large Density = \frac{Mass}{Volume} $$

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I got 137.04 with that formula

OpenStudy (perl):

checking

OpenStudy (perl):

$$\Large { Density = \frac{Mass}{Volume} \\~\\ \iff \\Volume = \frac{Mass}{Density } } $$

OpenStudy (perl):

and you are correct

OpenStudy (anonymous):

Thank you